A drug is ordered 20​ mg/kg/d divided into q6h doses IM.

The manufacturer recommends reconstituting 500 mg vial with 9.8 mL NS yielding 50​ mg/mL. How many milliliters will you administer for each dose if the client weighs 90​pounds?
Answer #1

90 pounds = 40.9 kgs can be rounded to 41 kgs

20 mg/ kg per day is 20 \times 41 = 820 mg in a day

Each dose is administered intramascularly every 6th hourly in a day = 820 \div 4 = 205 mg

Required medication = required dose \times available volume \div available dose

                                  = 205 \times 10 \div 500

                                  = 4.1

The client has prescribed 205 mg every 6th hourly and the nurse should administer 4.1 ml of the drug for each dose.